TWO DIMENSIONAL COMPACT SIMPLE RIEMANNIAN MANIFOLDS ARE BOUNDARY DISTANCE RIGID

LEONID PESTOV AND GUNTHER UHLMANN

Abstract. We prove that knowing the lengths of geodesics joining points of the boundary of a two-dimensional, compact, simple Riemannian manifold with boundary, we can determine uniquely the Riemannian metric up to the natural obstruction.